OpenStudy (klepez):

Simplify the radical expression 4 square root 2/ square root 3

OpenStudy (anonymous):

\[\frac{ 4\sqrt{2} }{ \sqrt{3} }\]

OpenStudy (klepez):

\[\frac{ 4\sqrt{2} }{ \sqrt{3} }\]

OpenStudy (anonymous):

if you ^2 the equation it will get rid of all the radicals. or you can multiply it all by the square root of 3 to get the denominator

OpenStudy (klepez):

\[4\sqrt{2}\]

OpenStudy (klepez):

\[\frac{ 4\sqrt{2} }{ 3 }\]

OpenStudy (klepez):

\[\frac{ 4\sqrt{6} }{ 3 }\]

OpenStudy (anonymous):

the top would have to be multiplied by the square root of 3 as well.

OpenStudy (klepez):

\[\frac{ 8 }{ 3 }\]

OpenStudy (anonymous):

hod did you get 8/3

OpenStudy (klepez):

those are the choices

OpenStudy (anonymous):

ohh! well then do it this way